OpenStudy (anonymous):

Evaluate: 3t(t+2)-3t^2 for t=19 . Lost.

OpenStudy (anonymous):

3t*t+3t*2-3t^2=3t^2+6t-3t^2=6t can you continue??

OpenStudy (anonymous):

I substituted 19 into t. Which came out like this: 3(19) (19+2)-3(19)^2

OpenStudy (anonymous):

I'm pretty confused from there. I know distributed property must take place but when?

OpenStudy (anonymous):

Well my book is telling me it's 114, but I want to know how to actually solve it. I live Florida so I have exams coming up next week.

OpenStudy (whpalmer4):

Now just collapse the parentheses: \[3(19)(19+2)-3(19)^2 = 3(19)(21)-3(19)^2 = 3*19*21-3*19*19\] Be sure you do all the multiplications before the subtraction! Easier, though, would be to look at what marsss wrote and observe that you only have to evaluate 6t = 6(19) :-)

OpenStudy (anonymous):

Thanks men.

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

BTW: check out my youtube channel (: http://www.youtube.com/user/roffajin?feature=watch

OpenStudy (anonymous):

@Rott ohh bro! you like rocket team,so am i :)

OpenStudy (anonymous):

Really ? Sweet dude